# Auto-Encoded Supervision for Perceptual Image Super-Resolution

### v1

@@ Abstract
This work tackles the fidelity objective in the perceptual super-resolution (SR). Specifically, we address the shortcomings of pixel-level $\mathcal{L}_{\text{p}}$ loss ($\mathcal{L}_{\text{pix}}$) in the GAN-based SR framework. Since $\mathcal{L}_{\text{pix}}$ is known to have a trade-off relationship against perceptual quality, prior methods often multiply a small scale factor or utilize low-pass filters. However, this work shows that these circumventions fail to address the fundamental factor that induces blurring. Accordingly, we focus on two points: 1) precisely discriminating the subcomponent of $\mathcal{L}_{\text{pix}}$ that contributes to blurring, and 2) only guiding based on the factor that is free from this trade-off relationship. We show that they can be achieved in a surprisingly simple manner, with an Auto-Encoder (AE) pretrained with $\mathcal{L}_{\text{pix}}$. Accordingly, we propose the Auto-Encoded Supervision for Optimal Penalization loss ($\mathcal{L}_{\text{AESOP}}$), a novel loss function that measures distance in the AE space  [^1] AE space indicates the space after the decoder, not the bottleneck., instead of the raw pixel space. By simply substituting $\mathcal{L}_{\text{pix}}$ with $\mathcal{L}_{\text{AESOP}}$, we can provide effective reconstruction guidance without compromising perceptual quality. Designed for simplicity, our method enables easy integration into existing SR frameworks. Experimental results verify that AESOP can lead to favorable results in the perceptual SR task.
@@ 2 Related work
Fidelity-oriented SR. The pioneering works [13, 25], CNN-based [39, 68, 9, 47, 45, 70, 16] and Swin Transformer [40] based methods [5, 34, 6, 63, 36] have shown remarkable improvements. The majority of these methods employ $\mathcal{L}_{\text{pix}}$ as their sole objective. Thus, they minimize the expected per-pixel distance, which leads them to estimate the average over multiple solutions [20, 41]. This leads to a high PSNR score but is empirically shown to be blurry. In our work, the optimal fidelity point is the optimal estimation of these fidelity-oriented SR methods.
@@ 3 Revisiting per-pixel loss in perceptual SR
with $\hat{y}$ as an estimator of $y$, and $\mu_{y}=\operatorname*{arg\,min}_{\mu}\mathbb{E}_{y}[\mathcal{L}(y,\mu)]$ and $\mu_{\hat{y}}=\operatorname*{arg\,min}_{\mu}\mathbb{E}_{y}[\mathcal{L}(\hat{y},\mu)]$. With Gaussian prior on $\mathcal{L}$, the two terms above are further simplified as follows:
@@
Overall, since VE is the sole factor inducing the PD trade-off, it is straightforward that we can safely reduce SE without harming perception. By minimizing SE for a given VE, we can obtain the maximum fidelity for a given perception level; the ideal PD trade-off  [^3] Note that zero SE alone does not indicate optimal perception.. At the same time, a sufficient level of VE should be preserved for visual quality. Accordingly, the following sections will elaborate on designing a novel loss that minimizes SE while preserving VE, taking a step toward the optimal perceptual SR network.
Revisiting prior methods. Observations above share some key concepts with the well-known perception-distortion (PD) trade-off [2]. However, we highlight aspects of $\mathcal{L}_{\text{pix}}$ that are often overlooked in most training methods, with the terms defined above. Specifically, previous approaches aim to avoid blurring by either 1) introducing LPF before loss calculation [69, 11] or 2) simply applying a small coefficient for $\mathcal{L}_{\text{pix}}$ [30, 54, 37]. Below, we will show that both are misinterpretating the blurring phenomenon.
@@
Meanwhile, applying a small coefficient to $\mathcal{L}_{\text{pix}}$ misguidedly treats all aspects of $\mathcal{L}_{\text{pix}}$ as contributing to blurring. This indiscriminative approach unintentionally weakens SE reduction. When combined with the adversarial loss, this also leads to redundant SE, failing to reach the optimal PD trade-off. This is because adversarial loss works in a task-blind, unsupervised manner: it improves realism but does not consider the alignment between the input image and the network output. As this significantly hinders SE convergence [37, 59], strong guidance on the SE factor is required to prevent high SE. However, the small coefficient greatly reduces this guidance, resulting in unnecessary fidelity loss and a suboptimal PD trade-off. Additionally, since $\mathcal{L}_{\text{pix}}$ fundamentally enforces VE reduction, it can also never achieve optimal perception despite applying a small scale factor. See Appendix.F for discussions and graphical illustrations.
@@ 4.1 Auto-Encoder pretraining
Designing the fidelity bias feature space. Our aim is to provide reconstruction supervision focused exclusively on SE. Given that the only components of SE are the fidelity biases ($\mu_{y},\mu_{\hat{y}}$), our task simplifies into estimating an operator $\psi$ that estimates the fidelity bias of a given image as ${\psi(\cdot):=\operatorname*{arg\,min}_{\mu}\mathbb{E}[\mathcal{L}(\cdot,\mu)]}$. To maintain simplicity, we employ a basic Auto-Encoder (AE) to construct a differentiable approximation of this operator. This architectural choice is grounded by the nature of SR, where the HR images are conditioned by the LR images and the scale factor. Thus, we model this relationship as $y\sim p(y|x)=p(y|\phi(y,s))$, where $x$ is the observed LR image, $\phi$ is the ×$\times s$ downsampling function. In essence, HR images are derived from a distribution conditioned on themselves and the scale factor $s$. Additionally, the definition of $\psi$ involves minimizing the expected loss over this conditional distribution. Thus, our method is followed by pretraining the AE to reconstruct the input with the $\mathcal{L}_{p}$ loss. Accordingly, the pretrained AE will act as a differentiable approximation of $\psi$, which can decompose the fidelity bias of images and can also be directly plugged into the training framework.
@@
AE pretraining. To design a fidelity bias estimator, we pretrain our AE to approximate ${\psi(\cdot):=\operatorname*{arg\,min}_{\mu}\mathbb{E}[\mathcal{L}(\cdot,\mu)]}$ for $y\sim p(y|x)=p(y|\phi(y,s))$. Thus, the AE is pretrained to condition itself by estimating $\phi(y,s)$ and then reconstructs $y$. That is, it learns to consecutively estimate the LR and HR image by minimizing $\mathcal{L}_{\text{p}}$ for each as follows:
@@ 6 Conclusion
Appendix A Implementation and experimental details
Network architecture and weight initialization.
@@ Appendix C Additional results for the bicubic SR task
AESOP on recent backbone network architecture. We report additional quantitative results on the benchmark datasets in Tab.8. First, we employ DRCT [18], a recent state-of-the-art method that enhances the network architecture within the fidelity-oriented SR framework. We reproduce LDL on top of DRCT, and compare it with $\mathcal{L}_{\text{AESOP}}$. AESOP is shown to be effective in these recent network architectures as well, outperforming the baseline in both fidelity and perceptual quality.
@@
Additional perception-distortion trade-off curves. We provide extensive visualizations of the perception-distortion trade-off curves in Fig.12-18. For CALGAN [48], we present only a single data point rather than the full perception-distortion trade-off curve, as its official weights are not publicly available. Extensive results show that AESOP leads to substantial performance improvements against baselines in terms of the perception-distortion trade-off. Aligning to Tab.8, AESOP also often outperforms CALGAN even without MoE-discriminator proposed in CALGAN. Additionally, we observe that AESOP often results in larger improvements for Swin Transformer-based methods (e.g., SwinIR, DRCT) compared to CNN-based methods (e.g., RRDB). This is likely because these models have greater capacity and benefit more from the enhanced reconstruction guidance provided by AESOP. However, there are instances where AESOP does not always lead to improved performance. Specifically, AESOP often fails to enhance performance on the Manga109 [44] dataset, which is consistent with the unexpected trade-off behaviors observed across most methods in this dataset. This limitation arises because Manga109 consists predominantly of comic images, which typically lack the fine-grained textures found in photorealistic datasets. The absence of such textures poses a challenge for perceptual SR methods, including AESOP, which are specifically designed to enhance and preserve realistic textures. Consequently, without the presence of these detailed textures, AESOP’s advantages in minimizing fidelity bias and preserving perceptual variance are less pronounced, leading to suboptimal performance in this particular dataset.
@@
Table 8: Additional quantitative evaluation on benchmark datasets. We also provide quantitative results of CALGAN [48] and DRCT [18]. CALGAN is a recent work improving perceptual-oriented losses, while DRCT made improvements in the SR network architecture. AESOP mostly outperforms CALGAN [48] even without the MoE-discriminator. However, note that enhancements to network architectures and perceptual-oriented losses are beyond the scope of this work. The focus of this work is on the fidelity loss term $\mathcal{L}_{\text{pix}}$ under the perceptual SR framework. The best results of each group are highlighted in bold. Additionally, refer to the PD trade-off curve in Fig.12-18.
Metric Method Set14 Mg109 Gen100 Urb100 DIV2K B100 LSDIR
AE-PSNR ESRGAN [54] 30.280 31.165 32.663 27.198 31.668 28.991 27.636
SPSR [42] 30.602 31.351 32.670 27.508 31.737 29.029 27.881
LDL [37] 31.180 32.608 33.823 28.488 32.597 29.595 28.625
AESOP (Ours) 31.341 32.843 33.956 28.529 32.740 29.737 28.812
LR-PSNR ESRGAN [54] 43.892 43.908 45.259 42.879 45.689 43.823 42.718
SPSR [42] 43.835 44.359 44.656 42.666 44.717 42.719 42.364
LDL [37] 46.497 47.603 48.184 45.975 47.793 45.307 45.295
AESOP (Ours) 46.625 48.188 48.653 46.280 48.272 45.837 45.571
Table 9: AE-PSNR and LR-PSNR scores on the RRDB-backbone.
Appendix D Additional visualization
Qualitative examples on benchmark datasets. To further illustrate the effectiveness of our method, we present an additional qualitative comparison between AESOP against the baseline method LDL [37] on the bicubic $\times$4 SR task. We provide results of tested methods, AESOP and LDL, on top of the SwinIR backbone (Fig.20 and Fig.21) and the RRDB backbone (Fig.22 and Fig.23). As can be seen, the proposed AESOP significantly improves the perceptual quality while effectively suppressing visual artifacts observed in the baseline method.
Figure 8: Loss map comparison between high-pass filtered (HPF) loss and $\mathcal{L}_{\text{AESOP}}$. Refer to Appendix.E for further discussion.
Appendix E Further discussion on AESOP
@@
Figure 9: Loss scale comparison between $\mathcal{L}_{\text{pix}}$ and $\mathcal{L}_{\text{AESOP}}$, before and after applying the loss coefficients of each. The loss trajectory of $\mathcal{L}_{\text{pix}}$ before applying the coefficient (green) is visualized by scaling the original loss (red). The loss trajectory of $\mathcal{L}_{\text{AESOP}}$ is visualized as-is, since we do not scale it. Refer to Appendix.E for further discussion.
Intuitions on $\mathcal{L}_{\text{AESOP}}$ based on loss scales. In Fig.9, we compare the loss scales of $\mathcal{L}_{\text{pix}}$ and $\mathcal{L}_{\text{AESOP}}$, both before and after applying their loss coefficients. Before the loss coefficients are applied, $\mathcal{L}_{\text{pix}}$ (green) exhibits greater loss values than $\mathcal{L}_{\text{AESOP}}$ (blue). This observation aligns with our theoretical analysis and construction of the Auto-Encoder, where $\mathcal{L}_{\text{AESOP}}$ only penalizes a subcomponent of $\mathcal{L}_{\text{pix}}$. Specifically, while $\mathcal{L}_{\text{pix}}$ minimizes both perceptual variance (VE) and fidelity bias induced error (SE), our carefully designed $\mathcal{L}_{\text{AESOP}}$ only targets the SE term, leading to lower loss values. Consequently, the gap between the green loss trajectory and the blue one quantifies the VE loss component embedded within $\mathcal{L}_{\text{pix}}$. When the loss coefficients are applied, $\mathcal{L}_{\text{AESOP}}$ (blue) provides an order of magnitude stronger reconstruction guidance compared to scaled $\mathcal{L}_{\text{pix}}$ (red). Regardless of this increased fidelity guidance, SR networks trained with $\mathcal{L}_{\text{AESOP}}$ do not have to suffer from blurring and can achieve improved perceptual quality over $\mathcal{L}_{\text{pix}}$.

### v2

@@ Abstract
This work tackles the fidelity objective in the perceptual super-resolution (SR) task. Specifically, we address the shortcomings of pixel-level $\mathcal{L}_{\text{p}}$ loss ($\mathcal{L}_{\text{pix}}$) in the GAN-based SR framework. Since $\mathcal{L}_{\text{pix}}$ is known to have a trade-off relationship against perceptual quality, prior methods often multiply a small scale factor or utilize low-pass filters. However, this work shows that these circumventions fail to address the fundamental factor that induces blurring. Accordingly, we focus on two points: 1) precisely discriminating the subcomponent of $\mathcal{L}_{\text{pix}}$ that contributes to blurring, and 2) only guiding based on the factor that is free from this trade-off relationship. We show that they can be achieved in a surprisingly simple manner, with an Auto-Encoder (AE) pretrained with $\mathcal{L}_{\text{pix}}$. Accordingly, we propose the Auto-Encoded Supervision for Optimal Penalization loss ($\mathcal{L}_{\text{AESOP}}$), a novel loss function that measures distance in the AE space  [^1] AE space indicates the space after the decoder, not the bottleneck., instead of the raw pixel space. By simply substituting $\mathcal{L}_{\text{pix}}$ with $\mathcal{L}_{\text{AESOP}}$, we can provide effective reconstruction guidance without compromising perceptual quality. Designed for simplicity, our method enables easy integration into existing SR frameworks. Extensive experiments demonstrate the effectiveness of AESOP.
@@ 2 Related work
Fidelity-oriented SR. The pioneering works [13, 25], CNN-based [39, 68, 9, 47, 45, 70, 16] and Swin Transformer [40] based methods [5, 34, 6, 63, 36] have shown remarkable improvements. The majority of these methods employ the per-pixel loss $\mathcal{L}_{\text{pix}}$ as their sole objective. This leads them to estimate the average over multiple solutions [20, 41], resulting in high PSNR scores but is empirically shown to be blurry. In our work, the optimal fidelity point is the optimal estimation of these fidelity-oriented SR methods.
@@ 3 Revisiting per-pixel loss in perceptual SR
with $\hat{y}$ as an estimator of $y$, and $\mu_{y}=\operatorname*{arg\,min}_{\mu}\mathbb{E}_{y}[\mathcal{L}(y,\mu)]$ and $\mu_{\hat{y}}=\operatorname*{arg\,min}_{\mu}\mathbb{E}_{\hat{y}}[\mathcal{L}(\hat{y},\mu)]$. For $\mathcal{L}_{\text{2}}$, the two terms above are further simplified as follows:
@@
Overall, since VE is the sole factor inducing the PD trade-off, it is straightforward that we can safely reduce SE without harming perception. By minimizing SE for a given VE, we can obtain the maximum fidelity for a given perception level; the ideal PD trade-off  [^3] Zero SE alone does not indicate optimal perception.. At the same time, a sufficient level of VE should be preserved for visual quality. Accordingly, the following sections will elaborate on designing a novel loss that minimizes SE while preserving VE, taking a step toward the optimal perceptual SR network.
Revisiting prior methods. Observations above share some key concepts with the well-known perception-distortion (PD) trade-off [2]. However, we highlight aspects of $\mathcal{L}_{\text{pix}}$ that are often overlooked in most training methods, with the terms defined above. Specifically, previous approaches aim to avoid blurring by either 1) introducing LPF before loss calculation [69, 11] or 2) simply applying a small coefficient for $\mathcal{L}_{\text{pix}}$ [30, 54, 37]. Below, we will show that both are misinterpreting the blurring phenomenon.
@@
Meanwhile, applying a small coefficient to $\mathcal{L}_{\text{pix}}$ misguidedly treats all aspects of $\mathcal{L}_{\text{pix}}$ as contributing to blurring. This indiscriminative approach unintentionally weakens SE reduction. When combined with the adversarial loss, this also leads to redundant SE, failing to reach the optimal PD trade-off. This is because adversarial loss works in a task-blind, unsupervised manner: it improves realism but does not consider the alignment between the input image and the network output. As this significantly hinders SE convergence [37, 59], strong guidance on the SE factor is required to prevent high SE. However, the small coefficient greatly reduces this guidance, resulting in unnecessary fidelity loss and a suboptimal PD trade-off. Additionally, since $\mathcal{L}_{\text{pix}}$ fundamentally enforces VE reduction, it can also never achieve optimal perception despite applying a small scale factor. See Appendix.E for discussions and graphical illustrations.
@@ 4.1 Auto-Encoder pretraining
Designing the fidelity bias feature space. Our aim is to provide reconstruction supervision focused exclusively on SE. Given that the only components of SE are the fidelity biases ($\mu_{y},\mu_{\hat{y}}$), our task simplifies into estimating an operator $\psi$ that estimates the fidelity bias of a given image as ${\psi(\cdot):=\operatorname*{arg\,min}_{\mu}\mathbb{E}[\mathcal{L}(\cdot,\mu)]}$. To maintain simplicity, we employ a basic Auto-Encoder (AE) to construct a differentiable approximation of this operator. This architectural choice is grounded by the nature of SR, where the HR images are conditioned by the LR images and the scale factor. Thus, we model this relationship as $y\sim p(y|x)$ with $x\equiv\phi(y,s)$, where $\phi$ is the ×$\times s$ downsampling function. Additionally, the definition of $\psi$ involves minimizing the expected loss over this conditional distribution. Thus, we pretrain the AE with $\mathcal{L}_{p}$ to learn the forward mapping $x\leftarrow y$, and consecutively, the inverse mapping $y\leftarrow x$. Now, the pretrained AE will act as a differentiable approximation of $\psi$, which can decompose the fidelity bias of images and can also be directly plugged into the training framework.
@@
AE pretraining. To design a fidelity bias estimator, we pretrain our AE to approximate ${\psi(\cdot):=\operatorname*{arg\,min}_{\mu}\mathbb{E}[\mathcal{L}(\cdot,\mu)]}$ for $y\sim p(y|x)$ with $x\equiv\phi(y,s)$. Thus, the AE consecutively estimates the low-resolution counterpart $x$ and then reconstructs $y$. Accordingly, the pretraining objective is straightforward as follows:
@@ 6 Conclusion
Acknowledgement
This work was supported in part by MSIT/IITP (No. 2022-0-00680, 2020-0-01821, RS-2019-II190421, RS-2024-00459618, RS-2024-00360227, RS-2024-00437102, RS-2024-00437633), and MSIT/NRF (No. RS-2024-00357729).
@@ Acknowledgement
Appendix A Implementation and experimental details
Network architecture and weight initialization.
Appendix B Evaluation on real-world SR datasets
@@ Appendix C Additional results for the Bicubic SR task
AESOP on recent backbone network architecture. We report additional quantitative results on the benchmark datasets in Tab.9. First, we employ DRCT [18], a recent state-of-the-art Swin Transformer-based method that leverages dense residual connections within a fidelity-oriented SR framework. We implement LDL on top of DRCT and compare it to our proposed AESOP. AESOP consistently outperforms the baseline in terms of both fidelity and perceptual quality, demonstrating its effectiveness even with advanced network architectures. Notably, the performance improvement is more significant compared to the RRDB backbone, suggesting that AESOP may yield even greater benefits with larger-capacity network architectures.
@@
Metric Method Set14 Manga109 General100 Urban100 DIV2K-val BSD100 LSDIR
AE-PSNR ESRGAN [54] 30.280 31.165 32.663 27.198 31.668 28.991 27.636
SPSR [42] 30.602 31.351 32.670 27.508 31.737 29.029 27.881
LDL [37] 31.180 32.608 33.823 28.488 32.597 29.595 28.625
AESOP (Ours) 31.341 32.843 33.956 28.529 32.740 29.737 28.812
LR-PSNR ESRGAN [54] 43.892 43.908 45.259 42.879 45.689 43.823 42.718
SPSR [42] 43.835 44.359 44.656 42.666 44.717 42.719 42.364
LDL [37] 46.497 47.603 48.184 45.975 47.793 45.307 45.295
AESOP (Ours) 46.625 48.188 48.653 46.280 48.272 45.837 45.571
Table 8: Quantitative comparison between the proposed method (AESOP) and baseline methods. We report AE-PSNR and LR-PSNR scores using the RRDB backbone. AE-PSNR measures how accurately the method estimates fidelity bias factors, while LR-PSNR evaluates how well the generated images align with the input LR image. The best result in each group is highlighted in bold.
@@
Table 9: Additional quantitative evaluation on benchmark datasets. We also provide quantitative results of CALGAN [48] and DRCT [18]. CALGAN is a recent work improving perceptual-oriented losses, while DRCT made improvements in the SR network architecture. AESOP mostly outperforms CALGAN [48] even without the MoE-discriminator. However, note that enhancements to network architectures and perceptual-oriented losses are beyond the scope of this work. The focus of this work is on the fidelity loss term $\mathcal{L}_{\text{pix}}$ under the perceptual SR framework. The best results of each group are highlighted in bold. Additionally, refer to the PD trade-off curve in Fig.14-20.
Appendix D Further discussion on AESOP
@@
Figure 8: Loss map comparison between high-pass filtered (HPF) loss and $\mathcal{L}_{\text{AESOP}}$. Refer to Appendix.D for further discussion.
Figure 9: The loss trajectory of $\mathcal{L}_{\text{pix}}$ before applying the coefficient (green) is visualized by scaling the original loss (red). The loss trajectory of $\mathcal{L}_{\text{AESOP}}$ (blue) is visualized as-is, since we do not scale it. Refer to Appendix.D for further discussion.
Figure 10: Comparison between feature inversion results obtained from deep features extracted by VGG and our proposed AE. Deep features of VGG lose important low-level features crucial for a reconstruction loss. Meanwhile, the carefully chosen network architecture and the pretraining objective of our AE enable precise control over which information to remove (blur-inducing high-frequency patterns) and preserve (structural edges).
Intuitions on $\mathcal{L}_{\text{AESOP}}$ based on loss scales. In Fig.9, we compare the loss scales of $\mathcal{L}_{\text{pix}}$ and $\mathcal{L}_{\text{AESOP}}$, both before and after applying their loss coefficients. Before the loss coefficients are applied, $\mathcal{L}_{\text{pix}}$ (green) exhibits greater loss values than $\mathcal{L}_{\text{AESOP}}$ (blue). This observation aligns with our theoretical analysis and construction of the Auto-Encoder, where $\mathcal{L}_{\text{AESOP}}$ only penalizes a subcomponent of $\mathcal{L}_{\text{pix}}$. Specifically, while $\mathcal{L}_{\text{pix}}$ minimizes both perceptual variance (VE) and fidelity bias induced error (SE), our carefully designed $\mathcal{L}_{\text{AESOP}}$ only targets the SE term, leading to lower loss values. Consequently, the gap between the green loss trajectory and the blue one quantifies the VE loss component embedded within $\mathcal{L}_{\text{pix}}$. After the loss coefficients are applied to each reconstruction loss, $\mathcal{L}_{\text{AESOP}}$ (blue) provides an order of magnitude stronger reconstruction guidance compared to scaled $\mathcal{L}_{\text{pix}}$ (red). Regardless of this strengthened fidelity guidance, SR networks trained with $\mathcal{L}_{\text{AESOP}}$ do not have to suffer from blurring and can achieve improved perceptual quality over $\mathcal{L}_{\text{pix}}$.
@@
Comparison between $\mathcal{L}_{\text{percep}}$ and $\mathcal{L}_{\text{AESOP}}$. The proposed loss $\mathcal{L}_{\text{AESOP}}$ and the perceptual loss $\mathcal{L}_{\text{percep}}$ share the characteristic of utilizing a pretrained neural network for guidance. However, they differ fundamentally in their objectives and mechanisms. Below, we clarify these differences in two different aspects.
First, the primary objectives of these two losses differ significantly.$\mathcal{L}_{\text{percep}}$ belongs to the category of perceptual-oriented losses. Its main purpose is to explicitly improve perceptual quality by providing high-level semantics and textural guidance. Accordingly, $\mathcal{L}_{\text{percep}}$ measures the discrepancy between the SR and HR images within a high-dimensional feature space derived from a pretrained feature extractor (such as VGG [51]), where the high-dimensional space captures additional semantic and textural details beyond those available in the raw pixel domain. In contrast, $\mathcal{L}_{\text{AESOP}}$ is fundamentally a reconstruction (fidelity) loss that provides guidance based on low-level features, similar to the conventional $\mathcal{L}_{\text{pix}}$, but specifically tailored for perceptual SR tasks so that it does not show conflicts with perceptual-oriented losses. $\mathcal{L}_{\text{AESOP}}$ employs an Auto-Encoder (AE) architecture with a low-dimensional bottleneck, pretrained for low-level reconstruction. Due to its design and pretraining objective, the AE inherently compresses the input and selectively discards certain information, while preserving important low-level features. Consequently, the Auto-Encoded output contains less information compared to the original image, as opposed to the enriched, high-dimensional features used in $\mathcal{L}_{\text{percep}}$.
Second, the underlying mechanism and the information each feature encoder embeds are different. In order to utilize a feature encoder as a loss function in low-level vision tasks, precise control over which information to remove and preserve is important. Considering that a reconstruction loss in perceptual SR task should (1) provide sufficient reconstruction guidance while (2) avoid blurring; feature encoders should be able to preserve important low-level features while removing blur-inducing factors. However, feature encoders pretrained on image classification tasks (such as VGG) naturally discard many low-level features not relevant to classification, resulting in uncontrollable loss of critical reconstruction information. In contrast, the carefully designed AE preserves essential low-level features, particularly structural edges, while the blur inducing perceptual variance factors are removed.
We empirically verify these properties through feature inversion results shown in Fig.10. Clearly, deep features extracted from VGG omit critical low-level reconstruction details. On the other hand, our AE-derived deep features successfully retain sharp edges and structural alignment while the blur inducing high-frequency textural information is removed as intended. Overall, this verifies that $\mathcal{L}_{\text{percep}}$ cannot act as a standalone reconstruction loss in low-level vision tasks, while $\mathcal{L}_{\text{AESOP}}$ can, and is even shown to outperform conventional $\mathcal{L}_{\text{pix}}$ through extensive experiments.
Disclaimer.
We clarify that the improvement in perceptual scores by raising $\mathcal{L}_{\text{AESOP}}$ is since it does not hinder the perceptual-oriented guidance provided by perceptual-oriented losses under the SRGAN-framework. $\mathcal{L}_{\text{AESOP}}$ itself will not guide towards realism. We keep improvements in perceptual-oriented losses out of the scope of this work.
Appendix E Further intuition regarding the PD trade-off